TAG Metallic Inc. deals in main products, Metallic Doors and Window Frames.  There have been recent concerns about profit fluctuations, and there are concerns whether the products are appropriately costed and priced. TAG Metallic uses a cost margin pricing policy and uses a pre-determined overhead absorption rate based on direct labour hours for doors and machine hours for windows. The rate is calculated at the start of the year based on the following information: -

Production overhead                                       GHS 1,500,000

Direct labour hours                                                     60,000

Machine hours                                                             100,000

Set up hours                                                                2,000

The following additional information is provided upon preliminary investigations:


Metallic Doors             Window Frames

Production Volume   250 units                        100 units

Direct materials                                   GHS 350                     GHS 450

Ending value of materials                    GHS 80                       GHS 120

Net value of Work in progress GHS 60                       GHS (80)

Direct labour hours                                   20                                 25

Machine hours                                           50                                 60

Set up hours                                              2                                   2.5

Margin                                                      2/7                              20%

Direct labour rate                                 GHS 20                         GHS 15


Further analysis of production overheads and supporting information reveals the following:-



Set up

Maintenance

Cutting

Assembly

Total


GHS

GHS

GHS

GHS

GHS

Indirect

60000

60000

100000

100000

320000

Materials






Supervision



80000

120000

200000

Support costs

60000

60000

135000

180000

435000

Depreciation


25000

150000

50000

225000


You have been asked to analyse the information and make relevant calculations using traditional and activity costing methods, to inform decisions on pricing.


NB: maintenance and cutting are based on machine hours whiles assembly is on labour hours.



Required:


i. Calculate the predetermined overhead absorption rates used by TAG Metallic at the start of the year.                                                           

[4 marks]

ii. Calculate the standard cost and the sales price of Doors and Windows, using the predetermined overhead absorption rate.  

iii.  Calculate an activity based overhead absorption rate  

iv. Calculate the cost and the sales price of Doors and Windows using the activity based overhead absorption rate.

B). Explain the differences in the operations of Coca-Cola bottling Ghana and Kpogas Furniture  

C) Distinguish between Prime cost and Conversion cost, and could there be a situation where the two distinct cost terms are the same in cedi value? Give reason for your stand.
Answer #1

Production overhead                                       GHS 1,500,000

Direct labor hours                                                     60,000

pre-determined overhead absorption rate based on direct labor-hours for doors = 1500000/ 60000 = 25

Machine hours                                                             100,000

the predetermined overhead absorption rate based on machine hours for windows. = 1500000/100000 = 15

Set up hours                                                                2,000

Calculate the standard cost and the sales price of Doors and Windows, using the predetermined overhead absorption rate.

                                                           Metallic Doors              Window Frames

Production Volume 250 units                             100 units

Total Direct materials Utilised                        87500                                45000

Ending value of materials                     20000                          12000

The net value of Work in progress           15000                      (8000)

Direct labour                                             100000                              37500

Total Overhead                               125000                              90000

Total standard cost                                                  347500                                176500

(using the predetermined overhead absorption rate )

Margin                                                                     99286                                   35300

Total Sales Value                                                     446786                                  211800

Sales price                                                             1787.14                                      2118

                                                                                                                                                                     

analysis of production overheads and supporting information reveals the following:-


Set up

Maintenance

Cutting

Assembly

Total


GHS

GHS

GHS

GHS

GHS

Indirect Materials

60000

60000

100000

100000

320000

Supervision



80000

120000

200000

Support costs

60000

60000

135000

180000

435000

Depreciation


25000

150000

50000

225000

Total1200001450004650004500001180000
Total Production Overhead



1500000
Miscellaneous Overhead (B.F.)



320000

Maintenance & Cutting Overhead = 145000 + 465000 = 610000 are based on machine hours

Recovery Rate = 610000/100000 = 61

Assembly Overhead = 450000 are based on labour hours

Recovery Rate = 450000/ 60000 = 7.5

Set up Overhead = 120000 are based on Set up hours     

Recovery Rate = 120000/2000 = 60

Miscellaneous Overhead = 320000 Assumed to be based on Set up hours

Recovery Rate = 320000/2000 = 160

Calculate the cost and the sales price of Doors and Windows using the activity based overhead absorption rate.

                                                           Metallic Doors              Window Frames

Direct labour hours                                    20                                   25

Machine hours                                              50                                   60

Set up hours                                                2                                   2.5

Production Volume    250 units                             100 units

Total Direct materials Utilised                           87500                                45000

Ending value of materials                        20000                          12000

Net value of Work in progress           15000                      (8000)

Direct labour                                       100000                              37500

Maintenance & Cutting Overhead                          762500                                 366000

Assembly Overhead                                                  37500                                 18750

Set up Overhead                                                       30000                                15000

Mscellaneous Overhead 80000 40000

Total standard cost                                                 1132500                            526250

(using the activity based overhead absorption rate )

Margin                                                                       323571                           105250

Total Sales Value                                                   1456071                          631500

Sales price                                                            5824.28                               6315


answered by: Dr. R